OpenStudy (anonymous):

ok i don't even know where to begin with this problem. "a certain right triangle contains one leg whose length is 7 cm. The length of the other leg and the hypotenuse are consecutive integers. Find these missing lengths."

OpenStudy (anonymous):

See if one leg measures n cm, then the hypotenuse is n+1 cm now by pythagoras theorem 7^2 + n^2 = (n+1)^2

OpenStudy (amistre64):

7^2 + c = (c+1)^2

OpenStudy (amistre64):

is openstudy lagging today? or is it just me?

OpenStudy (anonymous):

no its driving everyone CRAZY

OpenStudy (anonymous):

IM FED UP OF IT...ARRRRRRRGGGGGGGGGHHHHHHHHH

OpenStudy (amistre64):

him banner; turns all green and fuzzy when he gets angry lol

OpenStudy (anonymous):

Tiffany... continuing 49 =2n+1 n=24

OpenStudy (amistre64):

at least we got some cute pics to bid the time with ;)

OpenStudy (anonymous):

why dont we tell someone amistre..this is crazy..totally agree on the pics part...;)

OpenStudy (anonymous):

so the hypot. would be 25?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol and thank you. :)

OpenStudy (amistre64):